# Latent Factor Analysis in Short Panels

### v2

@@ E.5.3 Feasible CLT for the FA estimators
Assumption 3 holds e.g. for conditionally homoskedastic mds, and for ARCH processes (see below). Let $\mathscr{Z}:=V_{\varepsilon}^{-1/2}ZV_{\varepsilon}^{-1/2}$. Then, using Lemma 1, under Assumptions 2 and 3, we have $V[\mathscr{Z}_{t,t}]=\psi(0)+2\kappa$, $V[\mathscr{Z}_{t,s}]=\psi(t-s)+q+\kappa$ and $Cov(\mathscr{Z}_{t,t},\mathscr{Z}_{s,s})=\psi(t-s)$, where $\psi(t-s):=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i}Cov(w_{i,t}^{2},w_{i,s}^{2})\sigma_{ii}^{2}$. Quantity $\psi(t-s)$ depends on the difference $t-s$ only, by stationarity. The other covariance terms between elements of $\mathscr{Z}$ vanish. Then, we have $\Omega=[\psi(0)-2q]D(0)+\sum_{h=1}^{T-1}\psi(h)D(h)+(q+\kappa)I_{T(T+1)/2}$, where $D(0)=\sum_{t=1}^{T}vech(E_{t,t})vech(E_{t,t})^{\prime}$ and $D(h)=\tilde{D}(h)+\bar{D}(h)$ with $\tilde{D}(h)=\sum_{t=1}^{T-h}[vech(E_{t,t})vech(E_{t+h,t+h})^{\prime}+vech(E_{t+h,t+h})vech(E_{t,t})^{\prime}]$ and $\bar{D}(h)=\sum_{t=1}^{T-h}vech(E_{t,t+h}+E_{t+h,t})vech(E_{t,t+h}+E_{t+h,t})^{\prime}$ for $h=1,...,T-1$, and where $E_{t,s}$ denote the $T\times T$ matrix with entry $1$ in position $(t,s)$ and 00 elsewhere. Hence, with $Z=V_{\varepsilon}^{1/2}\mathscr{Z}V_{\varepsilon}^{1/2}$, we get a parametrization $\Omega_{Z}(V_{\varepsilon},\vartheta)$ for $V[vech(Z)]$ with $\vartheta=(q+\kappa,\psi(0)-2q,\psi(1),...,\psi(T-1))^{\prime}$.
@@
Under Assumptions 1-3 and A.1-A.6, we have:
$M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}=\sum_{h=1}^{T-1}[\psi(h)+q+\kappa]M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\bar{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}.$ (E.24)
Hence, the parametric structure $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}(V_{\varepsilon},G,\tilde{\vartheta})$ depends linearly on vector $\tilde{\vartheta}$ that stacks the $T-1$ parameters $\psi(h)+q+\kappa$, for $h=1,...,T-1$. It does not involve parameter $\psi(0)$, i.e., the quartic moment of errors, because the asymptotic expansion of the LR statistic does not involve the diagonal terms of $Z$. Moreover, the unknown parameters appear through the linear combinations $\psi(h)+q+\kappa$ that are the scaled variances of the out-of-diagonal elements of $Z$. We can estimate the unknown parameters in $\tilde{\vartheta}$ by least squares applied on (E.24), using the nonparametric estimator $M_{\hat{\boldsymbol{X}}}\hat{\Omega}_{Z^{*}}M_{\hat{\boldsymbol{X}}}$ defined in Proposition 2, after half-vectorization and replacing $V_{\varepsilon}$ and $G$ by their FA estimates. It yields a consistent estimator of $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}$ incorporating the restrictions implied by Assumption 3.
@@
Under Assumptions 1-3 and A.1-A.6, and $\sqrt{n}\sum_{m=1}^{J_{n}}B_{m,n}^{2}=o(1)$, up to pre- and post-multiplication by an orthogonal matrix and its transpose, we have $\hat{\Omega}_{Z^{*}}=\boldsymbol{R}^{\prime}\tilde{\Xi}_{n}\boldsymbol{R}+o_{p}(1)$, where $\tilde{\Xi}_{n}=[\psi_{n}(0)-2q_{n}]D(0)+\sum_{h=1}^{T-1}\psi_{n}(h)D(h)+(q_{n}+\kappa_{n})I_{T(T+1)/2}+(q_{n}+\xi_{n})vech(I_{T})vech(I_{T})^{\prime}$ and $\xi_{n}:=\frac{1}{n}\sum_{m=1}^{J_{n}}\sum_{i\neq j\in I_{m}}\sigma_{ii}\sigma_{jj}$.
@@
$\boldsymbol{R}^{\prime}\tilde{\Xi}_{n}\boldsymbol{R}$ $=$ $[\psi_{n}(0)+q_{n}+3\kappa_{n}]\boldsymbol{R}^{\prime}D(0)\boldsymbol{R}+\sum_{h=1}^{T-1}[\psi_{n}(h)+q_{n}+\kappa_{n}]\boldsymbol{R}^{\prime}D(h)\boldsymbol{R}$ (E.25)
@@
depends on $T+1$ linear combinations of the elements of $\vartheta_{n}=(q_{n}+\kappa_{n},\psi_{n}(0)-2q_{n},\psi_{n}(1),...,\break\psi_{n}(T-1))^{\prime}$ and $\xi_{n}-\kappa_{n}$. Thus, the linear system (E.25) is rank-deficient to identify $\vartheta_{n}$. Moreover, in Assumption A.3 (b), $\kappa_{n}$ is defined as a double sum over squared covariances scaled by $n$, and is assumed to converge to a constant $\kappa$. Such a convergence is difficult to assume for $\xi_{n}$ since $\xi_{n}$ is a double sum over products of two variances scaled by $n$.
@@
We apply half-vectorization on (E.25), replace the LHS by its consistent estimate $\hat{\Xi}$, and plug-in the FA estimates in the RHS. From Lemma 9, least squares estimation on such a linear regression yields consistent estimates of linear combinations $\psi(0)+q+3\kappa$ and $\psi(h)+q+\kappa$ for $h=1,\dots,T-1$. Consistency of those parameters applies independently of $\xi_{n}-\kappa_{n}$ converging as $n\rightarrow\infty$, or not. [^40] To see this, write the half-vectorization of the RHS of (E.25) as $\chi\eta_{n}$, where $\chi$ is the $\frac{p(p+1)}{2}\times(T+1)$ matrix of regressors and $\eta_{n}$ the $(T+1)\times 1$ vector of unknown parameters. Then, $vech(\hat{\Omega}_{Z^{*}})=\hat{\chi}\eta_{n}+o_{p}(1)$, by Lemma 9, the consistency of the FA estimates, and the last column of $\chi$ not depending on unknown parameters. Thus, $\hat{\eta}_{n}:=(\hat{\chi}^{\prime}\hat{\chi})^{-1}\hat{\chi}^{\prime}vech(\hat{\Omega}_{Z^{*}})=\eta_{n}+o_{p}(1)$. In particular, we also have $\widehat{\xi_{n}-\kappa_{n}}=\xi_{n}-\kappa_{n}+o_{p}(1)$. In order to identify the components of $\vartheta$, we need an additional condition. We use the assumption $\psi(T-1)=0$. That condition is implied by serial uncorrelation in the squared standardized errors after lag $T-1$, that is empirically relevant in our application with monthly returns data. Then, parameter $q+\kappa$ is estimated by $\widehat{\psi_{n}(T-1)+q_{n}+\kappa_{n}}$, and by difference we get the estimators of $\psi(0)-2q$ and $\psi(h)$, for $h=1,...,T-2$.
@@
Let us now discuss the case of ARCH errors. Suppose the $w_{i,t}$ follow independent ARCH(1) processes with Gaussian innovations that are independent across assets, i.e., $w_{i,t}=h_{i,t}^{1/2}z_{i,t},\ z_{i,t}\sim IIN(0,1),\ h_{i,t}=c_{i}+\alpha_{i}w_{i,t-1}^{2}$ with $c_{i}=1-\alpha_{i}$. Then $E[w_{i,t}]=0$, $E[w_{i,t}^{2}]=1$, $\eta_{i}:=V[w_{i,t}^{2}]=\frac{2}{1-3\alpha_{i}^{2}}$, $Cov(w_{i,t}^{2},w_{i,t-h}^{2})=\eta_{i}\alpha_{i}^{h}$. Moreover, $E[w_{i,t}w_{i,r}w_{i,s}w_{i,p}]=0$ if one index among $t,r,s,p$ is different from all the others. Indeed, without loss of generality, suppose $t$ is different from $s,p,r$. By the law of iterated expectation: $E[\varepsilon_{i,t}\varepsilon_{i,s}\varepsilon_{i,p}\varepsilon_{i,r}]=E[E[\varepsilon_{i,t}|\{z_{i,\tau}^{2}\}_{\tau=-\infty}^{\infty},\{z_{i,\tau}\}_{\tau\neq t}]\varepsilon_{i,s}\varepsilon_{i,p}\varepsilon_{i,r}]=E[h_{i,t}^{1/2}E[z_{i,t}|z_{i,t}^{2}]\varepsilon_{i,s}\varepsilon_{i,p}\varepsilon_{i,r}]=0$. Then, Assumption 3 holds. The explicit formula of $\Omega$ involves $\psi(h)=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}\frac{2\alpha_{i}^{h}}{1-3\alpha_{i}^{2}}\sigma_{ii}^{2}$, for $h=0,1,...,T-1$. Hence, setting $\psi(T-1)=0$ is a mild assumption for identification purpose since $\alpha_{i}^{T-1}$ is small. If $\alpha_{i}=0$ for all $i$, i.e., no ARCH effects, we have $\psi(0)=2q$ and $\psi(h)=0$ for $h>0$, so that $\Omega=(q+\kappa)I_{\frac{T(T+1)}{2}}$.

### v3

@@ B.5.3 Feasible CLT for the FA estimators
Assumption 11 holds e.g. for conditionally homoskedastic mds, and for ARCH processes (see below). The stationarity condition is used only to ensure that $Cov(w^{2}_{i,t},w^{2}_{i,s})$ depends on the difference tst-s rather than on $t$ and $s$ individually. This assumption could be replaced by this weaker condition without affecting the results. This condition allows us to significantly reduce the number of parameters to estimate, making estimation more tractable. Let $\mathscr{Z}:=V_{\varepsilon}^{-1/2}ZV_{\varepsilon}^{-1/2}$. Then, using Lemma 2, under Assumptions 2 and 11, we have $V[\mathscr{Z}_{t,t}]=\psi(0)+2\kappa$, $V[\mathscr{Z}_{t,s}]=\psi(t-s)+q+\kappa$ and $Cov(\mathscr{Z}_{t,t},\mathscr{Z}_{s,s})=\psi(t-s)$, where $\psi(t-s):=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i}Cov(w_{i,t}^{2},w_{i,s}^{2})\sigma_{ii}^{2}$. Quantity $\psi(t-s)$ depends on the difference tst-s only, by stationarity. The other covariance terms between elements of $\mathscr{Z}$ vanish. Then, we have $\Omega=[\psi(0)-2q]D(0)+\sum_{h=1}^{T-1}\psi(h)D(h)+(q+\kappa)I_{T(T+1)/2}$, where $D(0)=\sum_{t=1}^{T}vech(E_{t,t})vech(E_{t,t})^{\prime}$ and $D(h)=\tilde{D}(h)+\bar{D}(h)$ with $\tilde{D}(h)=\sum_{t=1}^{T-h}[vech(E_{t,t})vech(E_{t+h,t+h})^{\prime}+vech(E_{t+h,t+h})vech(E_{t,t})^{\prime}]$ and $\bar{D}(h)=\sum_{t=1}^{T-h}vech(E_{t,t+h}+E_{t+h,t})vech(E_{t,t+h}+E_{t+h,t})^{\prime}$ for h=1,,T1h=1,...,T-1, and where $E_{t,s}$ denote the $T\times T$ matrix with entry $1$ in position $(t,s)$ and 0 elsewhere. Hence, with $Z=V_{\varepsilon}^{1/2}\mathscr{Z}V_{\varepsilon}^{1/2}$, we get a parametrization $\Omega_{Z}(V_{\varepsilon},\vartheta)$ for $V[vech(Z)]$ with $\vartheta=(q+\kappa,\psi(0)-2q,\psi(1),...,\psi(T-1))^{\prime}$.
@@
Under Assumptions 1-8, we have:
$M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}=\sum_{h=1}^{T-1}[\psi(h)+q+\kappa]M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\bar{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}.$ (B.24)
[…]
@@
Under Assumptions 1-8, and $\sqrt{n}\sum_{m=1}^{J_{n}}B_{m,n}^{2}=o(1)$, up to pre- and post-multiplication by an orthogonal matrix and its transpose, we have $\hat{\Omega}_{Z^{*}}=\boldsymbol{R}^{\prime}\tilde{\Xi}_{n}\boldsymbol{R}+o_{p}(1)$, where $\tilde{\Xi}_{n}=[\psi_{n}(0)-2q_{n}]D(0)+\sum_{h=1}^{T-1}\psi_{n}(h)D(h)+(q_{n}+\kappa_{n})I_{T(T+1)/2}+(q_{n}+\xi_{n})vech(I_{T})vech(I_{T})^{\prime}$ and $\xi_{n}:=\frac{1}{n}\sum_{m=1}^{J_{n}}\sum_{i\neq j\in I_{m}}\sigma_{ii}\sigma_{jj}$.
@@
$\boldsymbol{R}^{\prime}\tilde{\Xi}_{n}\boldsymbol{R}$ $=$ $[\psi_{n}(0)+q_{n}+3\kappa_{n}]\boldsymbol{R}^{\prime}D(0)\boldsymbol{R}+\sum_{h=1}^{T-1}[\psi_{n}(h)+q_{n}+\kappa_{n}]\boldsymbol{R}^{\prime}D(h)\boldsymbol{R}$ (B.25)
@@
depends on $T+1$ linear combinations of the elements of $\vartheta_{n}=(q_{n}+\kappa_{n},\psi_{n}(0)-2q_{n},\psi_{n}(1),...,\penalty-10000\psi_{n}(T-1))^{\prime}$ and $\xi_{n}-\kappa_{n}$. Thus, the linear system (B.25) is rank-deficient to identify $\vartheta_{n}$. Moreover, in Assumption 5 (b), $\kappa_{n}$ is defined as a double sum over squared covariances scaled by $n$, and is assumed to converge to a constant $\kappa$. Such a convergence is difficult to assume for $\xi_{n}$ since $\xi_{n}$ is a double sum over products of two variances scaled by $n$.
@@
We apply half-vectorization on (B.25), replace the LHS by its consistent estimate $\hat{\Xi}$, and plug-in the FA estimates in the RHS. From Lemma 10, least squares estimation on such a linear regression yields consistent estimates of linear combinations $\psi(0)+q+3\kappa$ and $\psi(h)+q+\kappa$ for $h=1,\dots,T-1$. Consistency of those parameters applies independently of $\xi_{n}-\kappa_{n}$ converging as $n\rightarrow\infty$, or not. [^13] To see this, write the half-vectorization of the RHS of (B.25) as $\chi\eta_{n}$, where $\chi$ is the $\frac{p(p+1)}{2}\times(T+1)$ matrix of regressors and $\eta_{n}$ the $(T+1)\times 1$ vector of unknown parameters. Then, $vech(\hat{\Omega}_{Z^{*}})=\hat{\chi}\eta_{n}+o_{p}(1)$, by Lemma 10, the consistency of the FA estimates, and the last column of $\chi$ not depending on unknown parameters. Thus, $\hat{\eta}_{n}:=(\hat{\chi}^{\prime}\hat{\chi})^{-1}\hat{\chi}^{\prime}vech(\hat{\Omega}_{Z^{*}})=\eta_{n}+o_{p}(1)$. In particular, we also have $\widehat{\xi_{n}-\kappa_{n}}=\xi_{n}-\kappa_{n}+o_{p}(1)$. In order to identify the components of $\vartheta$, we need an additional condition. We use the assumption $\psi(T-1)=0$. That condition is implied by serial uncorrelation in the squared standardized errors after lag T1T-1, that is empirically relevant in our application with monthly returns data. Then, parameter $q+\kappa$ is estimated by $\widehat{\psi_{n}(T-1)+q_{n}+\kappa_{n}}$, and by difference we get the estimators of $\psi(0)-2q$ and $\psi(h)$, for h=1,,T2h=1,...,T-2.
@@
Let us now discuss the case of ARCH errors. Suppose the $w_{i,t}$ follow independent ARCH(1) processes with Gaussian innovations that are independent across assets, i.e., $w_{i,t}=h_{i,t}^{1/2}z_{i,t},\ z_{i,t}\sim IIN(0,1),\ h_{i,t}=c_{i}+\alpha_{i}w_{i,t-1}^{2}$ with $c_{i}=1-\alpha_{i}$. Then $E[w_{i,t}]=0$, $E[w_{i,t}^{2}]=1$, $\eta_{i}:=V[w_{i,t}^{2}]=\frac{2}{1-3\alpha_{i}^{2}}$, $Cov(w_{i,t}^{2},w_{i,t-h}^{2})=\eta_{i}\alpha_{i}^{h}$. Moreover, $E[w_{i,t}w_{i,r}w_{i,s}w_{i,p}]=0$ if one index among $t,r,s,p$ is different from all the others. Indeed, without loss of generality, suppose $t$ is different from $s,p,r$. By the law of iterated expectation: $E[\varepsilon_{i,t}\varepsilon_{i,s}\varepsilon_{i,p}\varepsilon_{i,r}]=E[E[\varepsilon_{i,t}|\{z_{i,\tau}^{2}\}_{\tau=-\infty}^{\infty},\{z_{i,\tau}\}_{\tau\neq t}]\varepsilon_{i,s}\varepsilon_{i,p}\varepsilon_{i,r}]=E[h_{i,t}^{1/2}E[z_{i,t}|z_{i,t}^{2}]\varepsilon_{i,s}\varepsilon_{i,p}\varepsilon_{i,r}]=0$. Then, Assumption 11 holds. The explicit formula of $\Omega$ involves $\psi(h)=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}\frac{2\alpha_{i}^{h}}{1-3\alpha_{i}^{2}}\sigma_{ii}^{2}$, for h=0,1,,T1h=0,1,...,T-1. Hence, setting $\psi(T-1)=0$ is a mild assumption for identification purpose since $\alpha_{i}^{T-1}$ is small. If $\alpha_{i}=0$ for all $i$, i.e., no ARCH effects, we have $\psi(0)=2q$ and $\psi(h)=0$ for $h>0$, so that $\Omega=(q+\kappa)I_{\frac{T(T+1)}{2}}$.